OpenStudy (anonymous):

Q'R'S'T' is a dilation of QRST. If the area of QRST is 6 square units and the area of Q'R'S'T is 15 square units, what scale factor was used to dilate QRST? Explain your answer.

OpenStudy (anonymous):

L1 * W1 = 6 L1 (s) *W1(s) = 15 L1 * W1 *s^2 = 15 6 s^2 = 15 s = sqrt(15/6)

OpenStudy (anonymous):

Why do you need to square root it?

OpenStudy (anonymous):

It's because my scale factor increases each sidelength. Look at an example: If I start with a 3x2 rectangle, the area is 6. Now suppose I dilate the rectangle with a factor of 3. The rectangle is now a 9x3 and the new area is 27. The area did more than triple.